# Compactifications of phylogenetic systems and electrical networks

### v2

@@ 5.3. Plabic Tilings and Ranges
When $n=4$, the four Ptolemy diagrams are $Pt_{4}=$[Uncaptioned image]. Thus we have $|\mathbf{\xi}(\mathbf{\Omega}_{4})|=2^{0}+2^{4}+2^{4}+2^{4}=49,$ as seen in Figure 3, inside the dashed line. For $n=1,\dots,6$ we have $|\mathbf{\xi}(\mathbf{\Omega}_{n})|=1,2,8,49,373,3196$ respectively.
The proof of Theorem 17 can now be easily used to to count the images of compactified electrical networks:
Corollary 22.
The enumeration of the images of compactified electrical networks yields
Expanding this formula gives Corollary 2. For example, when $n=4,$ we have the total:
$(1/5)(5(49)+20(8)+10(4)+30(2)+5(1))=102.$ A summary of early results is given by Table 1.
$n$ $|\mathbf{\Omega}_{n}|$ $|{\mathbf{\Omega}}_{n}|$ $|\mathbf{\Psi}_{n}|$ $|\mathbf{\Psi}_{n}|$ $|\mathbf{\xi}(\mathbf{\Omega}_{n})|$ $|\mathbf{\sigma}(\mathbf{\Omega}_{n})|=|\mathbf{\xi}({\mathbf{\Omega}}_{n})|$
1 1 1 1 1 1 1
2 2 3 2 3 2 3
3 8 15 8 15 8 15
4 52 105 64 117 49 102
5 464 945 1024 1565 373 839
OEIS [A111088] [A001147] [A006125] [A136654] [?] [?]
Table 1. Enumeration of cells in each CW complex.

### v3

@@ 7.2. Traveling Salesman Polytopes
Theorem 35 shows that the set of incidence vectors corresponding to the cyclic orders in $\mathcal{O}_{c}(N)$ make up a face of STSP($n$). The 24 vectors associated to $N$ in Figure 2 are the vertices of the face of STSP(8) with $f$-vector $(1,24,96,186,210,145,60,13,1)$. For $N$ in Figure 2, the minimum is located simultaneously at the 24 vertices of STSP(8) where it takes the value 383/28. Thus, we answer Question (4) in the affirmative: By minimizing a given $W$, we can list the full set of cyclic orders allowing a planar construction for a circuit that gives $W$ upon measurement. Note that this listing is accomplished without knowing the actual structure of $N$, only its boundary measurements.